OpenStudy (anonymous):

how do i plot a trigonometric function in maple 15? f(t) = sec phi*t /4 how do i plot a trigonometric function in maple 15? f(t) = sec phi*t /4 @Mathematics

OpenStudy (anonymous):

\[\sec =\pi*t/4\]

OpenStudy (anonymous):



OpenStudy (anonymous):

plot(sec(t*\pi/4) ,x) should work. use the "pi" from the symbol palette

OpenStudy (anonymous):

if you don't want to see the asymptotes as actual lines, try plot( sec(pi*t/4), x, discont = true)